A 5-μC positive point charge is located at the origin and a 6 μC positive point charge is located at x = 0.00 m, y = 1.0 m. Find the coordinates of the point where the net electric field strength due to these charges is zero.

a) x = 0.00 m, y = 0.62 m

b) x = 0.00 m, y = 0.48 m

c) x = 0.00 m, y = 0.34 m

d) x = 0.00 m, y = 0.77 m
Answer #1

The point where electric field is zero will be on y axiz

Let that coordinate be (0,y)

Electric field at this point due to charge at origin=9*10^9*5*10^-6/y²

Electric field at this point due to the other charge=9*10^9*6*10^-6/(1-y)²

For net electric field to be zero, the electric field should be equal and opposite

9*10^9*5*10^-6/y²=9*10^9*5*10^-6/(1-y)²

5/y²=6/(1-y)²

y=0.48 m

Option b is correct